The bond has a coupon rate of 6.65 percent, it makes semiannual payments, and there are 2 months to the next coupon payment. A clean price of $1,031 and the par value is $1,000. What is the invoice price?
Answer #1

Bond's Invoice Price (Dirty Price) = Clean Price + Accumulated Interest

Annual Coupon (Interest) Rate = 6.65 %, Coupon Frequency: Semi-Annual, Par Value = $ 1000

Semi-Annual Coupon = C = Par Value x Annual Coupon Rate x 0.5 = 1000 x 0.0665 x 0.5 = $ 33.25

Time Between Two Coupons = t1 = 1 Semi-Annual Period = 6 months

Time to Next Coupon = t2 = 2 months

Accumulated Interest Period = t1 - t2 = 6 - 2 = 4 months

Accumulated Interest = (t/t1) x C = (4/6) x 33.25 = $ 22.167

Invoice Price = Clean Price + Accumulated Interest = 1031 + 22.167 = $ 1053.167 ~ $ 1053.17